OpenStudy (anonymous):

If all numbers are defined to be positive integers of the form 4n + 1 (i.e. 1,5,9,13,17, etc), show that the product of two of these numbers is also of the form 4n + 1.

OpenStudy (swissgirl):

=(4n+1)*(4n+1) =16n^2+4n+4n+1 =16n^2+8n+1 =4(4n^2+2n)+1 hmmm I wonder if this is correct

OpenStudy (swissgirl):

Hence its in the form 4k+1 where k=4n^2+2n

OpenStudy (anonymous):

That's where I was trying to go with this too, but it didn't seem to go anywhere.

OpenStudy (anonymous):

Seems like that's what should be done though; I just can't figure out how to bring it back around to 4n+1 without trivially dividing it by 4n+1.

OpenStudy (swissgirl):

Why u bringing it back to its original state???

OpenStudy (anonymous):

Because the point is to show that the product is of the same form...?

OpenStudy (swissgirl):

(4n+1)*4(n+1)=(4n+1)^2 =16n^2+4n+4n+1 =16n^2+8n+1 =4(4n^2+2n)+1 =4k+1 where k=4n^2+2n Hence we see that the product of of 4n+1 still retains the form

OpenStudy (anonymous):

Ah, yes! It was staring me in the face the entire time! Thanks!

OpenStudy (swissgirl):

I can tag someone else for their opinion if ud like :)

OpenStudy (swissgirl):

@dan815

ganeshie8 (ganeshie8):

this is exactly how we prove these... looks good to me !

OpenStudy (swissgirl):

Thanks :) I doubt myself

ganeshie8 (ganeshie8):

@ikram002p

OpenStudy (swissgirl):

oh nooo ikram is here <3

OpenStudy (ikram002p):

its correct Rachel :P

OpenStudy (swissgirl):

Thanks :D

OpenStudy (ikram002p):

are u working on mod ganesh ? or some other idea ?

OpenStudy (ikram002p):

btw rachel ur profile pic freaks me out xD

OpenStudy (swissgirl):

lmao Ik same here .... I gotta change it

ganeshie8 (ganeshie8):

no, the original proof is good enough :)

ganeshie8 (ganeshie8):

wait a second, i see a blunder

OpenStudy (swissgirl):

sh!t :(

ganeshie8 (ganeshie8):

you're assuming both numbers are same in ur proof, check once

OpenStudy (ikram002p):

its ok she can change and make it m anyway to see this in other way.(lazy mod) let 4m+1=1 mod 4 4n+1= 1 mod 4 (4n+1)(4m+1)= 1*1 mod 4 (4n+1)(4m+1)= 1 mod 4 (4n+1)(4m+1)= 4 k +1

OpenStudy (swissgirl):

Neat :)

ganeshie8 (ganeshie8):

(4n+1)*(4m+1) =4n(4m+1) + 1(4m+1) =4(4mn+n+m) + 1 =4K + 1

OpenStudy (ikram002p):

hehe

OpenStudy (swissgirl):

Yup I forgot abt that :) Thanks guys